Suppose you are given the following information about some hypothetical economy


and its national income accounts. Use this information to answer the questions that


follow. (Amounts are in millions of Kshs.)


Indirect Business Taxes 459.50


Corporate profits 604.45


Corporate profits taxes 234.70


Retained earnings 165.40


Proprietors' income 519.20


Rental Income 31.05


Net Interest 585.55


Exports 842.85


Imports 1190.20


Income Receipts from rest of world 427.8


Income Payments to rest of world 377.45


Net National Product 6190.4


Government expenditures for Goods and Services 1358.25


Personal current Transfer receipts 871.15


Social Security Taxes (employee & employer) 490.75


Personal Consumption expenditures 4892.85


Gross Private Domestic Investment 1082.45


Disposable personal income 5128.75


a. Find GDP (3 marks)


b. Find GNP (2 marks)


c. Find depreciation (capital consumption allowance) (2 marks)


d. Find National Income (3 marks)

1
Expert's answer
2021-11-02T10:21:53-0400

a

GDP = personal consumption expenditure + gross private domestic investment + government expenditure on goods and services + exports - imports

 

"= 4892.85 + 1082.45 + 1358.25 + 842.85 - 1190.20"

"GDP = 6986.2"


b

Depreciation = GNP - Net national product 

First we calculate GNP 

GNP = GDP + Income receipts from rest of world - Income payments to rest of world.

GNP = 6986.2 + 427.8 - 377.45

GNP = 7036.55


c

Now putting values in above depreciation formula.

Depreciation = 7036.55 - 6190.4

Depreciation = 846.15 


d

National income = Net national product - indirect business tax.

NNP = 6190.4

Indirect business tax = 459.50

Putting the value in formula.

National income = 6190.4 - 459.50

National income = 5730.9